OpenStudy (anonymous):

if sinx+cosx=sqrt(cosx) then prove that cosx-sinx=sqrt(sinx), this is from cbse class 11 problem

OpenStudy (anonymous):

Try squaring both sides first.

OpenStudy (anonymous):

(hint: this will only be true for certain values of x..)

OpenStudy (anonymous):

thanks i tried squaring on both sides but no help

OpenStudy (shubhamsrg):

=>multiply both sides by 1/sqrt(2) =>sin(x + pi/4) = sqrt(cos x /2) => sin^2 (x +pi/4) = cosx /2 =>2 sin^2 (x + pi/4) = cos x => 1- 2 sin^2(x+ pi/4) = 1 - cos x => cos ( 2x + pi/2) = 1- cosx => - cos( 2x) = 1- cos x => -2 cos^2 x +1 = 1 - cos x =>cos^2 x = cos x =>cos x = 0 , 1 both values dont satisfy original eqn,, where did i go wrong? o.O

OpenStudy (shubhamsrg):

wel,,cos x =1 satisfy original eqn, but not the latter one.. you sure the ques is right ? or please correct me in case i went wrong ?

OpenStudy (anonymous):

this question is correct only, it was given in the recent test for class 11 and i also have one more similar question that is: if tan square x = sinx * cosx then prove 2 - sin to the power of 6 x = 4 * sin square x = 3 * sin to the power of 4 x =1

OpenStudy (shubhamsrg):

ohh..lol// i will have 2 cos^2 x = cos x => cos x = 0 =>cos x = 1/2 when cos x =1/2, sin x = sqrt(3)/2 or - sqrt(3)/2 cos x = 0 isnt a solution,, maybe latter one satisfies it.. please confirm anyone ? and sorry for confusion above /

OpenStudy (shubhamsrg):

seems even latter one doesnt satisfy! :|

OpenStudy (anonymous):

Well there is no solution of x......I think we just need to prove from the given condition

OpenStudy (shubhamsrg):

if no x satisfies both,,how can it be true ?

OpenStudy (anonymous):

sinx + cosx = sqr(cosx) 1+2sin2x = cosx .....i cosx -sinx =sqr(sinx) 1-2sin2x = sinx....ii Now, from i and ii cosx + sinx =2 .......(which has no solution)

OpenStudy (shubhamsrg):

aha..cool.. conclusion will be given statement is false right ?

OpenStudy (raden):

but, (sinx+cosx)^2 = 1+sin2x not 1+2sin2x

OpenStudy (shubhamsrg):

@RadEn true !! but it wont effect the final eqn..

OpenStudy (raden):

sinx + cosx = sqrt(cosx) 1+sin2x = cosx sin2x = cosx - 1 ... i let k=cosx-sinx k^2 = 1-sin2x .... ii subs (i) to (ii) k^2 = 1-(cosx - 1) k^2 = 2-cosx k=sqrt(2-cosx) not prove! lol

OpenStudy (anonymous):

just a try can u use (a+b)^2 = (a-b)^2+4ab after squaring on both sides ...

OpenStudy (anonymous):

Not getting...

OpenStudy (anonymous):

@hartnn , And I am not typing...

OpenStudy (anonymous):

(sinx+cosx)^2 = cosx (cosx=sinx)^2 + 4sinx cosx = cosx (cosx=sinx)^2 = cosx - 2 2 sinx cosx (cosx=sinx)^2 = cosx -2 sin2x i dont know to reduce after this can anyone help.....

OpenStudy (anonymous):

\[\cos^2(x) - \sin^2(x) = \cos(2x)\] \[\implies (\cos(x) - \sin(x))(\cos(x) + \sin(x)) = \cos(2x)\] \[\cos(x) - \sin(x) = \frac{\cos(2x)}{\sqrt{\cos(x)}} = ??\]

OpenStudy (anonymous):

Prove RHS = \(\sqrt{sin(x)}\)